# Trajectory-based models in strong-field physics

### v1

@@ Abstract
keywords:
strong-field physics, semiclassical models, above-threshold ionization, generation of high-order harmonics, nonsequential double ionization, frustrated tunneling ionization
@@ 1 Introduction
Strong-field physics investigates the highly nonlinear phenomena that occur when intense laser pulses interact with atoms and molecules. Among the most well-known of these phenomena are above-threshold ionization (ATI), the formation of a high-energy plateau in the electron energy spectrum (high-order ATI), high-order harmonic generation (HHG), and nonsequential double ionization (NSDI) [see Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9] for reviews]. These complex phenomena require a robust theoretical description, and the theoretical approaches used in strong-field physics are continually being refined. These approaches are based on the direct numerical solution of the time-dependent Schrödinger equation (see, e.g., Refs. [10, 11, 12, 13]), the strong-field approximation (SFA) [14, 15, 16], and semiclassical models employing classical trajectories to describe electron motion in the continuum. Widely known examples of semiclassical models are the two-step model [17, 18, 19] and the three-step model [20, 21].
@@
Along with direct electrons, there are also electrons that are driven back to their parent ions by the oscillating laser field. These rescattered electrons can undergo large-angle scattering (up to $180^{\circ}$) on the parent ions. The rescattered electrons are responsible for the high-energy plateau observed in the ATI spectrum. The concept of rescattering has also allowed to understand the physical mechanism of the HHG and NSDI processes. In the case of HHG, the returning electron recombines with the parent ion, emitting a high-energy photon in the form of harmonic radiation. Alternatively, if the electron possesses sufficient kinetic energy upon return, it can liberate the second electron from the ion by, e.g., impact ionization or excitation with subsequent ionization by the laser field (see Ref. [6]). These processes are captured by the three-step model that advances the two-step model by incorporating the interaction of the photoelectron with the parent ion as its third and final step. Therefore, the three-step model provides a basis for understanding of rescattering-induced strong-field processes. Indeed, the three-step model has successfully explained a wide range of features observed in high-order ATI, HHG, and NSDI. Well-known examples include the cutoff energies in high-order ATI spectra [35] and HHG [20, 36], the maximum emission angles in the photoelectron angular distributions [37], and the characteristic momenta of recoil ions observed in NSDI [38, 39].
@@ 2 Trajectory-based models: Theoretical background
The electron trajectory in a semiclassical model is calculated from Newton’s equation of motion:
@@
Here, $\mathbf{r}$ is the electric field of the laser pulse, and $V\left(\mathbf{r},t\right)$ is the potential of the parent ion. For simplicity, we assume that the laser field is linearly polarized along the $z$-axis, $\mathbf{F}\left(t\right)=F_{z}\left(t\right)\mathbf{e}_{z}$, and we treat the tunneling problem in the quasistatic approximation (see, e.g., Ref [70]). In order to integrate Eq. (1), we need to know the initial conditions: the starting point and the initial velocity of the electron.
It should be stressed that the initial conditions are the weak point of the semiclassical approaches. Indeed, the initial conditions are to connect classical and quantum descriptions of strong-field processes, and therefore the choice of these conditions is a nontrivial task. In many trajectory-based models, it is assumed that the electron starts with zero initial velocity along the laser field $v_{0,z}=0$, but it can have a nonzero initial velocity $v_{0,\perp}$ in the perpendicular direction.
The simplest formula for the starting point (the tunnel exit point) $z_{e}$ can be derived by neglecting the ionic potential, i.e., considering a triangular potential barrier formed by the ground state energy and the laser field:
$r_{e}\left(t_{0}\right)=\frac{I_{p}}{F\left(t_{0}\right)}$ (2)
@@
If the interaction of the ionized electron with the parent ion can be modelled by the Coulomb potential, the tunnel exit point can be found by using the separation of the tunneling problem in parabolic coordinates. We define the parabolic coordinates as $\xi=r+z$, $\eta=r-z$, and $\phi=\arctan\left(y/x\right)$. The tunnel exit point $z_{e}=-\eta_{e}/2$ is then determined by the solution of the equation:
$-\frac{\beta_{2}\left(F\right)}{2\eta}+\frac{m^{2}-1}{8\eta^{2}}-\frac{F\eta}{8}=-\frac{I_{p}\left(F\right)}{4},$ (4)
@@
It should be mentioned that an effective potential for the outer electron that accounts for the laser field, the Coulomb potential, and the polarization effects of the ionic core was found in Refs. [71, 72, 73] by using the adiabatic approximation:
@@
$-\frac{\beta_{2}\left(F\right)}{2\eta}+\frac{m^{2}-1}{8\eta^{2}}-\frac{F\eta}{8}+\frac{\alpha_{I}F}{\eta^{2}}=-\frac{I_{p}\left(F\right)}{4},$ (8)
@@
Ionization times and the initial transverse velocities are often distributed by using the static tunneling rate (see, e.g., Ref. [70]):
@@
Modern semiclassical models apply various sampling strategies to distribute initial conditions and/or operate with classical trajectories in different ways. In order to illustrate these strategies, here we discuss semiclassical simulations of the ATI process. The simplest approach is to launch and propagate a large ensemble of classical trajectories. After the final momenta of all the trajectories are found, they can be distributed over the cells in momentum space, and thus the photoelectron momentum distribution is calculated. In strong-field physics, this approach is often referred to as the “shooting method", although it has nothing to do with a method for solving boundary value problems. Different sampling strategies can be used to implement the “shooting method." In the simplest approach, the initial conditions for classical trajectories are chosen randomly or from a uniform grid. In this case, the electron momentum distribution is obtained as:
@@
An alternative to the “shooting method" is provided by an approach based on solving the so-called “inverse problem", i.e., finding all trajectories leading to a given asymptotic momentum. The solution of the inverse problem allows for the avoidance of the need to calculate a large ensemble of classical trajectories. Furthermore, it controls the cusps and caustics that often emerge in classical simulations more efficiently than the “shooting method". However, the solution of the inverse problem is often very difficult. It is also clear that semiclassical models employing this approach are less versatile.
@@
As its name suggests, the classical trajectory Monte-Carlo method including tunneling (CTMC-T) is capable of accounting for the tunneling effect [82, 83]. In the CTMC-T method, the electron can tunnel through the potential barrier if it reaches the outer turning point. The latter is specified by the two following conditions: $p_{z}=0$ and $zF\left(t\right)<0$. The corresponding tunneling probability is calculated in accord with the WKB approximation (see Refs. [84, 85, 86]):
[…]
@@ 3 Semiclassical simulations of above-threshold ionization
Until about 10-15 years ago, semiclassical models had a significant drawback: they were not able to account for quantum interference effects. Indeed, it was not possible to reproduce interference patterns in electron momentum distributions within a purely classical treatment of electron dynamics after ionization. This deficiency has been overcome with the emergence of the trajectory-based Coulomb SFA (TCSFA) [44, 89], the quantum trajectory Monte Carlo model (QTMC) [90], the semiclassical two-step model (SCTS) [91], and the Coulomb quantum orbit strong-field approximation (CQSFA) [92, 93, 94, 95, 96] (the foundations of the CQSFA method are presented in Ref. [97]). All these semiclassical models are able to include interference effects in the two-step model with the ionic potential. The basic idea of all these approaches is to assign a certain phase to each trajectory. The contributions of different trajectories leading to the same final momentum are then added coherently.
The TCSFA extends the Coulomb-corrected strong-field approximation (CCSFA), see Refs. [98, 99]. In contrast to CCSFA, which considers the Coulomb field as a perturbation, the TCSFA approach treats the laser field and the atomic potential on equal footing. However, the TCSFA uses the first-order semiclassical perturbation theory (perturbation theory in action) [100] to account for the Coulomb potential in an interference phase assigned to every trajectory. The same first-order perturbation theory in action was also used in the QTMC model. The SCTS and the CQSFA account for the Coulomb potential beyond the semiclassical perturbation theory. While the SCTS model propagates large ensembles of classical trajectories, the CQSFA approach is based on the solution of the inverse problem. Since the TCSFA and CQSFA methods, although using classical trajectories, are derived from SFA and are closely related to this approach, we do not discuss them here. Undoubtedly, recent developments of both these methods deserve a separate thorough review article. In what follows, we consider the QTMC and the SCTS models.
@@
[…]
@@
[…]
@@
[…]
It is seen that the SCTS phase differs from that of the QTMC model [Eq. (15)]. Indeed, the Coulomb term enters with double weight in Eq. (16). This is due to the fact that the Coulomb potential is fully accounted for in the SCTS phase, whereas the QTMC phase includes the Coulomb effect within the semiclassical perturbation theory.
The SCTS model was compared with the direct numerical solution of the TDSE and predictions of the QTMC model [91, 69]. The electron momentum distributions calculated within all three approaches for the ionization of the hydrogen atom are shown in Fig. 1. It is seen that both semiclassical models reproduce the main important features of the TDSE results: the distributions are elongated along the polarization axis of the laser field, they have clear ATI rings, and a characteristic interference structure at small momenta. Both models underestimate the spread of the distribution along the polarization direction due to the initial condition $v_{z}=0$ (see Ref. [91]).
Figure 1: Two-dimensional momentum distributions for the H atom ionized by a laser pulse with a duration of $8$ cycles, wavelength of $800$ nm, and peak intensity of $0.9\times 10^{14}$ W/cm2 obtained from (a) the QTMC model, (b) solution of the TDSE, and (c) the SCTS model. The distributions are normalized to the total ionization yield. A logarithmic color scale in arbitrary units is used. The laser field is linearly polarized along the $z$-axis.
However, remarkable deviations are found in the low-energy part of the distributions. It is seen that for $\left|k\right|<0.3$ a.u., i.e., photoelectron energies below $U_{p}=0.2$ a.u., the 2D momentum distributions have fanlike interference structures. These interference structures are similar to those characteristic of Ramsauer-Townsend diffraction oscillations [106, 107, 108, 109]. The SCTS model reproduces the nodal pattern predicted by the TDSE, whereas the QTMC predicts fewer nodal lines, see Fig. 2. This is due to the underestimation of the Coulomb effect in the phase of the QTMC model, see Eq. (15).
Figure 2: Magnification of the distributions shown in Fig. 1 for $\left|k_{z}\right|$, $\left|k_{z}\right|<0.3$ a.u.
@@
Figure 3: Electron energy spectra calculated for the H atom ionized by a laser pulse with an intensity of $0.9\times 10^{14}$ W/cm2 and duration of $8$ cycles using numerical solution of the TDSE (thick light orange curve), the QTMC model (dotted green curve), and the SCTS model (solid red curve). Panels (a), (b), and (c) correspond to the wavelengths equal to 800 nm, 1000 nm, and 1200 nm, respectively. The energy spectra are normalized to the peak value. The dashed vertical lines show the energy equal to ponderomotive potential in panels (a) and (b).
When comparing semiclassical simulations with the direct numerical solution of the TDSE, we assume the latter to be exact. However, both the exact numerical solution of the TDSE and the obtaining of electron momentum distributions from the time-dependent wavefunction are nontrivial problems. Some methods used for extracting momentum distributions can lead to wrong results, see Ref. [110] for details.
@@
Currently, there is only a relatively small number of studies in which the entire high-energy above-threshold ionization (ATI) plateau up to an energy of 10Up has been calculated by launching an ensemble of classical trajectories, see, e.g., Refs. [113, 114], with the study of Ref. [114] pertaining to the ionization of a nanotip. Both these calculations are performed without considering quantum interference. Even though the momentum distributions of photoelectrons in the plateau region can be obtained by solving the TDSE or applying the quantitative rescattering theory, calculating these distributions using trajectory-based models is of interest. It should be noted that we are not aware of any work in which the momentum distributions of electrons in the plateau, their energy spectra, or angular distributions are calculated using any semiclassical model that accounts for the ionic potential and interference. This seems to be related to the fact that the bin size in momentum space, necessary for reliably describing interference oscillations, decreases with the increase in photoelectron energy and becomes small near the plateau cutoff. Indeed, the phase of semiclassical models contains a term proportional to the final electron energy: $\Phi\sim p^{2}/2$. Thus, when the final momentum of the photoelectron changes by $\Delta p$, the phase changes by $p\Delta p$. Comparing this change in phase with a phase difference of $\pi$ yields an estimate of the bin size in momentum space $\Delta p\sim\pi/p$. Obviously, decreasing the bin sizes requires a greater number of trajectories falling into it, i.e., a greater number of trajectories in the ensemble. This increases the computational costs of semiclassical modeling considerably. A possible solution would be the use of bins of variable size in momentum space and/or more sophisticated strategies for sampling initial conditions (e.g., a denser grid in regions of the $\left(t_{0},v_{0}\right)$-plane that predominantly yield rescattered trajectories, followed by multiplying the corresponding contributions by certain weights). Despite several efforts in this direction, the problem of accurately describing the high-energy ATI plateau within the framework of semiclassical modelling, considering interference and the potential of the ionic core, has not been solved yet.

### v2

@@ Abstract
: Laser Phys.
  • October 2025
Keywords: strong-field physics, semiclassical models, above-threshold ionization, generation of high-order harmonics, nonsequential double ionization, frustrated tunneling ionization
@@ 1 Introduction
Strong-field physics explores highly nonlinear processes that arise when atoms and molecules are exposed to intense laser fields. Among the most extensively studied effects are above-threshold ionization (ATI), the emergence of a high-energy plateau in the photoelectron spectrum (high-order ATI), high-order harmonic generation (HHG), and nonsequential double ionization (NSDI) [see Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9] for reviews]. These complex phenomena require a robust theoretical description, and the theoretical approaches used in strong-field physics are continually being refined. These approaches are based on the direct numerical solution of the time-dependent Schrödinger equation (see, e.g., Refs. [10, 11, 12, 13]), the strong-field approximation (SFA) [14, 15, 16], and semiclassical models employing classical trajectories to describe electron motion in the continuum. Widely known examples of semiclassical models are the two-step model [17, 18, 19] and the three-step model [20, 21].
@@
Along with direct electrons, there are also electrons that are driven back to their parent ions by the oscillating laser field. These rescattered electrons can undergo large-angle scattering (up to $180^{\circ}$) on the parent ions. The rescattered electrons are responsible for the high-energy plateau observed in the ATI spectrum. The concept of rescattering has also allowed to understand the physical mechanism of the HHG and NSDI processes. In the case of HHG, the returning electron recombines with the parent ion, emitting a high-energy photon in the form of harmonic radiation. Alternatively, if the electron possesses sufficient kinetic energy upon return, it can liberate the second electron from the ion by, e.g., impact ionization or excitation with subsequent ionization by the laser field (see Ref. [6]). These processes are captured by the three-step model that advances the two-step model by incorporating the interaction of the photoelectron with the parent ion as its third and final step, see Fig. 1. Therefore, the three-step model provides a basis for understanding of rescattering-induced strong-field processes. Indeed, the three-step model has successfully explained a wide range of features observed in high-order ATI, HHG, and NSDI. Well-known examples include the cutoff energies in high-order ATI spectra [35] and HHG [20, 36], the maximum emission angles in the photoelectron angular distributions [37], and the characteristic momenta of recoil ions observed in NSDI [38, 39].
Figure 1: The three-step model of strong-field ionization involves (i) electron tunneling through the effective potential barrier, (ii) acceleration by the laser field, and (iii) recollision with the parent ion, which can lead to high-order above-threshold ionization (via elastic scattering), high-harmonic generation (via recombination), or nonsequential double ionization (via the liberation of a second electron). The effective potential barrier (solid blue curve) results from the superposition of the Coulomb potential (dashed green curve) and the laser field potential (solid red curve).
@@ 2 Trajectory-based models: Theoretical background
In semiclassical models, the motion of an electron is determined by integrating Newton’s equation:
@@
In this notation, $\mathbf{r}$ denotes the electron position, while $V\left(\mathbf{r},t\right)$ represents the parent ion potential. For simplicity, we assume that the laser field is linearly polarized along the $z$-axis, $\mathbf{F}\left(t\right)=F_{z}\left(t\right)\mathbf{e}_{z}$, and we treat the tunneling problem in the quasistatic approximation (see, e.g., Ref [70]). In order to integrate Eq. (1), we need to know the initial conditions: the starting point and the initial velocity of the electron.
The choice of initial conditions is a delicate aspect of semiclassical models. The initial conditions serve to bridge the classical trajectory picture with the underlying quantum description, making their selection nontrivial. In many trajectory-based models, it is assumed that the electron starts with zero initial velocity along the laser field $v_{0,z}=0$, but it can have a nonzero initial velocity $v_{0,\perp}$ in the perpendicular direction.
A basic estimate for the tunnel exit point $z_{e}$ is obtained by neglecting the ionic potential, assuming a triangular barrier defined by the ground-state energy and the laser field:
$r_{e}\left(t_{0}\right)=\frac{I_{p}}{F\left(t_{0}\right)}.$ (2)
@@
When the Coulomb potential of the parent ion is included, the tunnel exit point can be obtained via separation in parabolic coordinates. We define the parabolic coordinates as $\xi=r+z$, $\eta=r-z$, and $\phi=\arctan\left(y/x\right)$. The tunnel exit point $z_{e}=-\eta_{e}/2$ is then determined by the solution of the equation:
$\frac{\beta_{2}\left(F\right)}{2\eta}+\frac{m^{2}-1}{8\eta^{2}}-\frac{F\eta}{8}=-\frac{I_{p}\left(F\right)}{4},$ (4)
@@
An effective potential including the laser field, Coulomb potential, and ionic core polarization has been introduced in Refs. [71, 72, 73] by using the adiabatic approximation:
@@
$\frac{\beta_{2}\left(F\right)}{2\eta}+\frac{m^{2}-1}{8\eta^{2}}-\frac{F\eta}{8}+\frac{\alpha_{I}F}{\eta^{2}}=-\frac{I_{p}\left(F\right)}{4},$ (8)
@@
A common approach assigns ionization times and initial transverse velocities according to the static tunneling rate (see, e.g., Ref. [70]):
@@
Modern semiclassical models apply various sampling strategies to distribute initial conditions and/or operate with classical trajectories in different ways. In order to illustrate these strategies, here we discuss semiclassical simulations of the ATI process. The simplest approach is to launch and propagate a large ensemble of classical trajectories. After the final momenta of all the trajectories are found, they can be distributed over the cells in momentum space, and thus the photoelectron momentum distribution is calculated. In strong-field physics, this approach is often referred to as the “shooting method”, although it has nothing to do with a method for solving boundary value problems. Different sampling strategies can be used to implement the “shooting method.” In the simplest approach, the initial conditions for classical trajectories are chosen randomly or from a uniform grid. In this case, the electron momentum distribution is obtained as:
@@
An alternative to the “shooting method” is provided by an approach based on solving the so-called “inverse problem”, i.e., finding all trajectories leading to a given asymptotic momentum. The solution of the inverse problem allows for the avoidance of the need to calculate a large ensemble of classical trajectories. Furthermore, it controls the cusps and caustics that often emerge in classical simulations more efficiently than the “shooting method”. However, the solution of the inverse problem is often very difficult. It is also clear that semiclassical models employing this approach are less versatile.
@@
The CTMC-T (classical trajectory Monte Carlo with tunneling) method explicitly includes tunneling events [82, 83]. In the CTMC-T method, the electron can tunnel through the potential barrier if it reaches the outer turning point. The latter is specified by the two following conditions: $p_{z}=0$ and $zF\left(t\right)<0$. The corresponding tunneling probability is calculated in accord with the WKB approximation (see Refs. [84, 85, 86]):
$w=\exp\left[-2\sqrt{2}\int_{z_{\mathrm{in}}}^{z_{\mathrm{out}}}\sqrt{V\left(z\right)-V\left(z_{\mathrm{in}}\right)}dz\right].$ (13)
@@ 3 Semiclassical simulations of above-threshold ionization
Until roughly a decade ago, semiclassical models could not reproduce quantum interference phenomena. Specifically, purely classical post-ionization dynamics failed to capture interference structures in electron momentum distributions. This deficiency has been overcome with the emergence of the trajectory-based Coulomb SFA (TCSFA) [44, 89], the quantum trajectory Monte Carlo model (QTMC) [90], the semiclassical two-step model (SCTS) [91], and the Coulomb quantum orbit strong-field approximation (CQSFA) [92, 93, 94, 95, 96] (the foundations of the CQSFA method are presented in Ref. [97]). In these approaches, each trajectory carries a phase, and the coherent sum of trajectories terminating at the same final momentum gives rise to interference patterns.
The TCSFA extends the Coulomb-corrected strong-field approximation (CCSFA), see Refs. [98, 99]. Unlike CCSFA, where the Coulomb potential is treated perturbatively, the TCSFA accounts for both laser and atomic potentials on an equal basis. However, the TCSFA uses the first-order semiclassical perturbation theory (perturbation theory in action) [100] to account for the Coulomb potential in an interference phase assigned to every trajectory. The same first-order perturbation theory in action was also used in the QTMC model. The SCTS and the CQSFA account for the Coulomb potential beyond the semiclassical perturbation theory. While the SCTS model propagates large ensembles of classical trajectories, the CQSFA approach is based on the solution of the inverse problem. Since the TCSFA and CQSFA methods, although using classical trajectories, are derived from SFA and are closely related to this approach, we do not discuss them here. Undoubtedly, recent developments of both these methods deserve a separate thorough review article. In what follows, we consider the QTMC and the SCTS models.
@@
$\Phi^{\textrm{QTMC}}\left(t_{0},\mathbf{v}_{0}\right)=-\mathbf{v}_{0}\cdot\mathbf{r}(t_{0})+I_{p}t_{0}-\int_{t_{0}}^{\infty}dt\,\left\{\frac{\mathbf{p}^{2}(t)}{2}-\frac{Z}{r\left(t\right)}\right\}.$ (15)
@@
$\Phi^{SCTS}\left(t_{0},\mathbf{v}_{0}\right)=-\mathbf{v}_{0}\cdot\mathbf{r}(t_{0})+I_{p}t_{0}-\int_{t_{0}}^{\infty}dt\,\left\{\dot{\mathbf{p}}(t)\cdot\mathbf{r}(t)+H[\mathbf{r}\left(t\right),\mathbf{p}(t)]\right\},$ (16)
@@
$\Phi^{\textrm{SCTS}}\left(t_{0},\mathbf{v}_{0}\right)=-\mathbf{v}_{0}\cdot\mathbf{r}(t_{0})+I_{p}t_{0}-\int_{t_{0}}^{\infty}dt\,\left\{\frac{\mathbf{p}^{2}(t)}{2}-\frac{2Z}{r\left(t\right)}\right\}.$ (17)
The SCTS phase differs from that of QTMC [Eq. (15)] because it incorporates the Coulomb potential beyond the perturbative approximation used in QTMC. Indeed, the Coulomb term enters with double weight in Eq. (16).
The SCTS model was compared with the direct numerical solution of the TDSE and predictions of the QTMC model [91, 69]. The electron momentum distributions calculated within all three approaches for the ionization of the hydrogen atom are shown in Fig. 2. Both semiclassical approaches capture the principal features of the TDSE distributions, such as elongation along the polarization axis and ATI rings, although they tend to underestimate the spread along the polarization direction due to the choice of $v_{z}=0$ (see Ref. [91]).
Figure 2: Two-dimensional momentum distributions for the H atom ionized by a laser pulse with a duration of 88 cycles, wavelength of $800$ nm, and peak intensity of $0.9\times 10^{14}$ W/cm2 obtained from (a,d) solution of the TDSE, (b,e) QTMC, and (c,f) the SCTS model. Panels (d), (e), and (f) show the maginifications for $\left|k_{z}\right|$, $\left|k_{z}\right|<0.3$ a.u. of the distributions presented in (a), (b), and (c), respectively. The distributions are normalized to the total ionization yield. A logarithmic color scale in arbitrary units is used. The laser field is linearly polarized along the $z$-axis.
However, remarkable deviations are found in the low-energy part of the distributions. It is seen that for $\left|k\right|<0.3$ a.u., i.e., photoelectron energies below $U_{p}=0.2$ a.u., the 2D momentum distributions have fanlike interference structures. These interference structures are similar to those characteristic of Ramsauer-Townsend diffraction oscillations [106, 107, 108, 109]. The SCTS model reproduces the nodal pattern predicted by the TDSE, whereas the QTMC predicts fewer nodal lines, see Fig. 2 (d)-(f). This is due to the underestimation of the Coulomb effect in the phase of the QTMC model, see Eq. (15).
@@
Figure 3: Electron energy spectra calculated for the H atom ionized by a laser pulse with an intensity of $0.9\times 10^{14}$ W/cm2 and duration of 88 cycles using numerical solution of the TDSE (thick light orange curve), the QTMC model (dotted green curve), and the SCTS model (solid red curve). Panels (a), (b), and (c) correspond to the wavelengths equal to 800 nm, 1000 nm, and 1200 nm, respectively. The energy spectra are normalized to the peak value. The dashed vertical lines show the energy equal to ponderomotive potential in panels (a) and (b).
For comparisons, TDSE results are taken as reference, though extracting momentum distributions from the wavefunction can be nontrivial and may introduce errors, see Ref. [110] for details.
@@
Currently, there is only a relatively small number of studies in which the entire high-energy above-threshold ionization (ATI) plateau up to an energy of 10Up has been calculated by launching an ensemble of classical trajectories, see, e.g., Refs. [113, 114], with the study of Ref. [114] pertaining to the ionization of a nanotip. Both these calculations are performed without considering quantum interference. Even though the momentum distributions of photoelectrons in the plateau region can be obtained by solving the TDSE or applying the quantitative rescattering theory, calculating these distributions using trajectory-based models is of interest. It should be noted that we are not aware of any work in which the momentum distributions of electrons in the plateau, their energy spectra, or angular distributions are calculated using any semiclassical model that accounts for the ionic potential and interference. This seems to be related to the fact that the bin size in momentum space, necessary for reliably describing interference oscillations, decreases with the increase in photoelectron energy and becomes small near the plateau cutoff. Indeed, the phase of semiclassical models contains a term proportional to the final electron energy: $\Phi\sim p^{2}/2$. Thus, when the final momentum of the photoelectron changes by $\Delta p$, the phase changes by $p\Delta p$. Comparing this change in phase with a phase difference of $\pi$ yields an estimate of the bin size in momentum space $\Delta p\sim\pi/p$. Obviously, decreasing the bin sizes requires a greater number of trajectories falling into it, i.e., a greater number of trajectories in the ensemble. This increases the computational costs of semiclassical modeling considerably. A possible solution would be the use of bins of variable size in momentum space and/or more sophisticated strategies for sampling initial conditions (e.g., a denser grid in regions of the $\left(t_{0},v_{0}\right)$-plane that predominantly yield rescattered trajectories, followed by multiplying the corresponding contributions by certain weights). Thus, achieving a reliable semiclassical description of the high-energy ATI plateau that includes both interference and the ionic potential remains an open challenge.